OpenStudy (anonymous):

help solving quadratic equation 8x+60=x^2+15x Help please please cx

Nnesha (nnesha):

set it equal to 0 so move 8x+60 to the right side

OpenStudy (anonymous):

so 0=x^2+15x- 8x+60?

Nnesha (nnesha):

nope 60 is positive at left side so subtract 60 both sides

OpenStudy (anonymous):

ohhh so 0=x^2+15x- 8x-60?

Nnesha (nnesha):

yes right now combine LIKE terms

OpenStudy (anonymous):

0=x^2+15x- 8x-60 ^ 7 0=x^2+7x-60

OpenStudy (anonymous):

@Nnesha what do i do next?

Nnesha (nnesha):

yes that's right now you can use quadratic formula to solve for x \[\huge\rm \frac{ {- }b \pm \sqrt{b^2 -4ac} }{ {2a} }\] or you can factor it \[ \large\rm \color{red}{1}x^2 +\color{Red}{ 7}x \color{reD}{- 60} = 0\]\[\large\rm \color{blue}{A}x^2 +\color{blue}{B}x+\color{blue}{C}=0\]

OpenStudy (anonymous):

the factoring one seems simpler lol

OpenStudy (anonymous):

i think thats what im suppose to do, is to factor it

Nnesha (nnesha):

alright to factor this one find two number if you multiply them you should get product of AC and if you add or subtract them you should get middle term b

Nnesha (nnesha):

A is leading coefficient C is constant term :-)

OpenStudy (anonymous):

-5 and 12

Nnesha (nnesha):

yep right so bec leading coefficient is just one GREAT! just write those two numbers with x (x + 1st number)(x+ 2nd number)

OpenStudy (anonymous):

(x+(-5))(x+12)

Nnesha (nnesha):

yes right + times -5 equal - so (x-5)(x+12)=0

Nnesha (nnesha):

if you have to solve for x set both parentheses equal to zero solve for x like this x-5 =0 and x+12 =0

OpenStudy (anonymous):

ohh ok, i would get 5 and -12

Nnesha (nnesha):

yep that's right!

OpenStudy (anonymous):

Thank you!, is that it?

OpenStudy (anonymous):

the answer

Nnesha (nnesha):

yep factors are (x-5)(x+12) solution or x value are 5 and -12

OpenStudy (anonymous):

Thank you so much! Cx

Nnesha (nnesha):

my pleasure :-)